OpenStudy (anonymous):

what is the simplified form of the following expression... 5/4 cubed

OpenStudy (anonymous):

(5/4) ^3 =(5*5*5)/(4*4*4)= 125/64

OpenStudy (anonymous):

Thanks! (: